OpenStudy (jamesreed30):

Please help will give medal!! I need to find the volume of a prism with sides of 2' 3' and 5'

OpenStudy (jango_in_dtown):

I guess it is a triangular prism.. Then the volume will be V=(1/2)* l*b*h=(1/2)*2*3*5=15

OpenStudy (jango_in_dtown):

@Jamesreed30

OpenStudy (jamesreed30):

It's not a triangular it's more of a rectangle one

OpenStudy (jamesreed30):

@jango_IN_DTOWN

OpenStudy (jango_in_dtown):

Then the formula is V=l*b*h=2*3*5=30

OpenStudy (jango_in_dtown):

@Jamesreed30

OpenStudy (jamesreed30):

What about the total area for the same figure?

OpenStudy (jamesreed30):

@jango_IN_DTOWN

OpenStudy (jango_in_dtown):

same as that of a cuboid. A= 2(lb+bh+hl)=62 (after calculation)

OpenStudy (jango_in_dtown):

@Jamesreed30

OpenStudy (jamesreed30):

And what about latereal area for the same figure

OpenStudy (jamesreed30):

@jango_IN_DTOWN

OpenStudy (jango_in_dtown):

2(lh + bh) = ???

OpenStudy (jango_in_dtown):

@Jamesreed30

OpenStudy (jamesreed30):

30? @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

depends on the picture... what is height, length, breadth we have to see

OpenStudy (jamesreed30):

The height is 5 length is 3

OpenStudy (jango_in_dtown):

2*5*(2+3)=6*5=30 correct

OpenStudy (jango_in_dtown):

@Jamesreed30

OpenStudy (jamesreed30):

Thank you! I have a few more questions if you wouldn't mind helping?

OpenStudy (jamesreed30):

@jango_IN_DTOWN

OpenStudy (jango_in_dtown):

If I can, I will. But post in a new question.. And if the question comes with a figure, then upload it

OpenStudy (jamesreed30):

I'm on my cell phone so I can't upload it or I would